( Discounted payback period) Gio's Restaurants is considering a project with the following expected cash flows. If the project's appropriate discount rate is 11 percent, what is the projects's discounted payback period?

expected cash flows:

Year Project Cash Flow
0 - $180 million
1 90 million
2 65 million
3 100 million
4 100 million

The Project's discounted payback period is ____
Answer #1

year discounted cash flow cumulative discounted cash flow

0 (180 m) (180 m)

1 90 m / (1.11) = 81.0811 ( 98.9189) [ 180 m -81.0811]

2 65 m / (1.11)2 = 52.7555 (43.8366) [ 180 m - 81.0811 -52.7555]

3 100 m / (1.11)3 = 73.1191 (26.9557) [ 180 m - 81.0811 - 52.7555 - 26.9557 ]

4 100 m / (1.11)4 = 65.8731 46.6654 [ 180 m - 81.0811 - 52.7555 - 26.9557- 65.8731]

discounted payback period = 3 + (26.9557 / 65.8731) =3.041 years